MArcos is 7 years older than his sister, Rosa. The sum of their ages is 1 less than 3 tomes Rosa's age. How old is each?
Found 2 solutions by ankor@dixie-net.com, iNeed:
Answer by ankor@dixie-net.com(22740)

Write an equation for each statement:
:
MArcos is 7 years older than his sister, Rosa.
M = R + 7
:
The sum of their ages is 1 less than 3 times Rosa's age.
M + R = 3R - 1
M = 3R - R - 1
M = 2R - 1
:
How old is each?
:
Substitute (R+7) for M in the above equation
R + 7 = 2R - 1
:
7 + 1 = 2R - R
:
R = 8 yrs is Rosa's age
then
M = 2(8) - 1 = 15 yrs is Marc's age
;
:
Check solution in the statement
The sum of their ages is 1 less than 3 times Rosa's age.
15 + 8 = 3(8) - 1

Answer by iNeed(3)

M=7+R.
M+R=1-3R.
7+R+R=3R-1.
7+2R=3R-1
8+2R=3R.
-2R -2R
8=R.
M=15.
8+15=24-1
23=23